A student fires a cannonball vertically upwards with a speed of 39m/s. Determine all unknowns and answer the following neglect drag and initial height and horizontal motion of the cannonball. What was the cannonball's maximum height? How long did the cannonball rise? What was the cannonball's total flight time?
1
Expert's answer
2020-09-29T09:40:23-0400

The cannonball's maximum height was


"h=\frac{v^2}{2g}=\frac{39^2}{2\cdot9.8}=77.6\text{ m}."

The time the cannonball rose was


"t=\sqrt{\frac{2h}{g}}=3.98\text{ s}."

The cannonball's total flight time was


"T=2t=7.96\text{ s}."